OpenStudy (anonymous):

Verify the identity. 1-sinx/cosx=cosx/1+sinx

OpenStudy (anonymous):

@mathslover any idea?

OpenStudy (arabpride):

I cant do trig >.< i terrible at it ~ sorry wish i cud help :/

OpenStudy (anonymous):

@arabpride know anyone that could?

OpenStudy (arabpride):

@*Insert Name Here*

OpenStudy (*insert name here*):

Multiply by the conjugate

OpenStudy (anonymous):

sorry but what part is the conjugate, (I'm really bad at pre calc)

OpenStudy (*insert name here*):

You have \(\Large \frac{1-sinx}{cosx}=\frac{cos}{1+sinx}\) Correct? Multiply by \(1+sinx\) \(\Large \frac{1-sin^2 x}{(cosx)(1+sinx)}\) Use the identity \(sin^2x+cos^2=1\) here So you end up with \(\Large \frac{cos^2x}{cosx(1+sinx)}\) Cancel a cos from the top and bottom leaving you with \(\Large \frac{cosx}{1+sinx}=\frac{cosx}{1+sinx}\)

OpenStudy (*insert name here*):

Make sense?

OpenStudy (anonymous):

yes perfect sense, THANK YOU!!

OpenStudy (*insert name here*):

You're welcome~